OpenStudy (yumeybitaa):

Can someone explain how to solve this type of ecuations ? (not the answer just how to do it ) Let f(x) = x2 - 8x + 5. Find f(-1).

OpenStudy (nirmalnema):

simply put (-1) in place of x in the given equation

OpenStudy (nirmalnema):

f(-1)=(-1)^2-8(-1)+5 =1+8+5 =14

OpenStudy (yumeybitaa):

so i just have to replace and solve ?

OpenStudy (nirmalnema):

yup..

OpenStudy (yumeybitaa):

and what happens if the ecuation is like this .... Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (yumeybitaa):

@nirmalnema

OpenStudy (nirmalnema):

first replace the x by 2 in the eq of g(x)=4-(6)(2)-7=-15 now the the value of g(x) in the function f(x) f(g(-15))=-15+8 =-7